OpenStudy (anonymous):

V=(1/3)pie r^2h solve for r

OpenStudy (anonymous):

that means that we want r by itself. So everything on the right hand side must go away except the r

OpenStudy (anonymous):

\[V=\frac{1}{3}\pi r^2h\] What do you want to get rid of first?

OpenStudy (anonymous):

um 1/3

OpenStudy (anonymous):

great that's division. Opposite of division is multiplication so we have to multiply both sides by 3\[3*V=3*\frac{1}{3}πr^2h\]\[3*V=πr^2h\]

OpenStudy (anonymous):

now we must get rid of pi and h so we have to divide both sides by pi and h....what do you get then?

OpenStudy (anonymous):

y by three

OpenStudy (anonymous):

can you write it again?

OpenStudy (anonymous):